# BioBERT-based Deep Learning and Merged ChemProt-DrugProt for Enhanced Biomedical Relation Extraction

### v1

@@ 2.4 Experimental details
In our experiments, we use $lr\_factor=0.0005$ and $warm\_up=1000$.
@@
For our reported results, CPR groups 7 and 8 were omitted due to their insufficient representation, which posed challenges for model generalization. Additionally, to augment the negative class (CPR:10), we introduced additional instances by selecting pairs of chemical and gene entities lacking any association in each input sentence. This approach contributed to the generation of more representative negative class instances.
@@ 2.5 Evaluation Metrics
We assessed the performance of our method using precision, recall, and $F_{1}$ score. Precision represents the ratio of correctly predicted mentions to the total set of predicted mentions for a specific entity, while recall signifies the ratio of correctly predicted mentions to the actual number of mentions. The $F_{1}$ score, calculated as the harmonic mean between precision and recall, provides a balanced measure of the model’s overall performance. To facilitate model evaluation, we randomly split our original merged training set into an 80/20 ratio, creating separate training and validation sets. The validation dataset was subsequently repurposed as the test set for final evaluation. During training, the 20% portion of the original merged training data served as the validation set for hyperparameter tuning and model optimization.
3 Results and Discussion
@@
The $F_{1}$ scores obtained from the test set across five individual runs are presented in Table 5. These results consistently demonstrate a pattern: BioBERT-Large consistently outperforms BioBERT-Base in terms of $F_{1}$ scores. Additionally, the inclusion of the top model leads to an enhancement in the $F_{1}$ score. However, it is important to note that this enhancement is not statistically significant, indicating that achieving state-of-the-art results does not necessarily require the adoption of a more complex system.
We conducted a comprehensive comparison between the performance of our merged dataset and the ChemProt dataset processed by Sun et al. [37] For this evaluation, we specifically focused on CPR groups 3, 4, 9, and 10, as they are common to both datasets processed by Sun et al. The substantial increase in the size of our dataset has led to a significant improvement in performance. The results for various CPR groups are presented in Table 6, with Precision, Recall and $F_{1}$ scores averaged over five execution runs. Notably, our RE model demonstrates a marked increase in performance across all positive groups (CPR 3, 4, and 9), attributed to the augmented sample counts within these groups. By concentrating on CPR groups 3, 4, and 9, shared between the ChemProt and DrugProt datasets, we aim to underscore the effectiveness of our approach. These groups serve as a crucial benchmark for evaluating the impact of our merged dataset, representing interactions consistently captured across both datasets. Analyzing the performance enhancements in these shared groups highlights how increased sample counts resulting from the merger of ChemProt and DrugProt contribute to the overall effectiveness of our relation extraction model. This focused comparison demonstrates the tangible benefits of our methodology in accurately predicting chemical-gene interactions.
ChemProt ChemProt-DrugProt
P R F P R F
CPR-3 0.7372 0.7407 0.7384 0.8690 0.8785 0.8735
CPR-4 0.7951 0.8236 0.8083 0.8894 0.9263 0.9073
CPR-9 0.6026 0.6976 0.6410 0.7620 0.8520 0.8040
CPR-10 0.9544 0.9415 0.9479 0.9725 0.9575 0.9649
Micro F1 0.0.7839 0.8875
Std dev 0.0025 0.0021
By focusing on CPR groups 3, 4, and 9, which are common between the ChemProt and DrugProt datasets, we aim to highlight the effectiveness of our approach. These specific groups serve as a crucial benchmark for evaluating the impact of our merged dataset, as they represent interactions that are consistently captured across both datasets. By analyzing the performance improvements in these shared groups, we can demonstrate how the increased sample counts resulting from the merging of ChemProt and DrugProt contribute to enhancing the overall effectiveness of our relation extraction model. This targeted comparison allows us to showcase the tangible benefits of our methodology in capturing and accurately predicting chemical-gene interactions that are pertinent to both datasets, thus affirming the utility and robustness of our approach.
4 Conclusion & Future Work
In conclusion, this study presents a robust methodology for enhancing relation extraction from biomedical texts, with a focus on chemical-gene interactions. By integrating the ChemProt and DrugProt datasets using a novel merging strategy, we demonstrated significant improvements in model performance, particularly in CPR groups shared between the datasets. Leveraging the BioBERT model and a fully connected network architecture, our approach effectively captures and predicts complex relationships within biomedical literature. The findings underscore the importance of dataset merging in augmenting sample counts and improving model accuracy, highlighting the potential of automated information extraction in biomedical research and clinical practice. Moving forward, further exploration of Natural Language Processing techniques in biomedicine holds promise for advancing knowledge discovery and innovation in the field.
In future work, several avenues for research and development present themselves to further enhance the capabilities of relation extraction in biomedical literature analysis. Firstly, investigating advanced methods for entity recognition and disambiguation could improve the accuracy of identifying chemical-gene interactions. Additionally, exploring techniques for incorporating contextual information and domain-specific knowledge into the model architecture may lead to more robust and context-aware relation extraction systems. Moreover, the integration of multi-modal data sources, such as images and structured data, could provide complementary information to further enrich relation extraction results. Furthermore, extending the analysis to encompass additional types of biomedical entities and relationships beyond chemicals and genes would broaden the scope of the research and facilitate a more comprehensive understanding of biological systems. Finally, collaboration with domain experts and stakeholders to validate and refine the extracted relations could enhance the practical utility of relation extraction systems in real-world biomedical applications. Overall, future work in this area holds the potential to drive significant advancements in biomedical research, clinical decision-making, and drug discovery efforts.
Acknowledgements

### v2

@@ 4.1 BioBERT Model Setup
In our experiments, we use lr_factor=0.0005lr\_factor=0.0005 and warm_up=1000warm\_up=1000.
@@
4.2 GCN-BERT Model Setup
With the effectiveness of our custom merged dataset established, we then proceeded to performing experiments with the GCN-BERT model to capture both local and global context for chemical-gene RE. Global context is important for incorporating information from co-occurring words. The GCN-BERT architecture is based on three key modules: vocabulary graph, Graph Convolutional Network and the combined GCN-BERT. The specifics of these modules are given in the sub-sections below.
4.2.1 Vocabulary Graph
The abstract text undergoes initial preprocessing, which includes the removal of stop words, punctuations, etc. The sentences are then tokenized using the BioBERT tokenizer [16]. The tokenizer maps these tokens to integers and handles out-of-vocabulary (OOV) words by splitting them into subwords and using ’##’ to represent split tokens. This processed input is used to create a vocabulary graph.
The number of nodes within this graph corresponds to the number of unique tokens in the entire corpus. Word nodes in the graph are represented with assigned integers. We calculate the weight of the edge between two word nodes using Pointwise Mutual Information (PMI). PMI is a metric that quantifies the probability of two words co-occurring. To compute the PMI score, we create a map of word-pair occurrences within a window of size 20. The frequency of these pairs across all windows is calculated to determine the PMI score between two word tokens. As a result, edges are established between word nodes when the PMI value is positive during the graph generation process. Therefore, if a particular chemical-gene pair appears multiple times throughout the corpus, that pair will have a positive PMI value.
The PMI formula [5] is expressed as follows:
$PMI(x,y)=\log\left(\frac{P(x,y)}{P(x)\cdot P(y)}\right)$ (1)
where:
$P(x,y)=\frac{\#SW(x,y)}{\#SW}\text{ and }P(x)=\frac{\#SW(x)}{\#SW}$ (2)
#SW(*) represents the number of sliding windows containing word pair or a particular word, and #SW represents the number of sliding windows.
In this context, the PMI values are normalized, resulting in Normalized PMI (NPMI) values within the range of [-1, 1]. A positive NPMI value indicates a significant semantic association between words, whereas a negative NPMI value suggests little to no semantic connection.
Figure 2: Architecture of GCN-BERT adapted from [24]
4.2.2 Graph Convolutional Network
Graph Convolutional Networks (GCNs) serve as a fundamental tool in analyzing complex chemical-gene interactions in the GCN-BERT model. The vocabulary graph is passed through a two-layer GCN, aiming to create graph embeddings that leverage the connections between neighboring nodes. GCNs are particularly valuable in the context of non-Euclidean data, as they excel at capturing intricate relations among entities.
One of the key strengths of GCNs lies in their ability to perform message passing within the graph. This means that the model can gather information from neighboring nodes, allowing it to grasp the broader context of chemical-gene interactions. The number of GCN layers used in our approach directly impacts the depth of information gathering, determining the maximum number of hops a node can make to capture information even from nodes that are not directly connected.
Our vocabulary graph is represented as G (N, E), where N is the set of nodes, with a node for each token in the corpus, and E denotes the set of edges within the graph. The vocabulary graph serves as a structure through which information about the semantic relations among words is propagated. The GCN model performs convolutions over this graph, aggregating information from the nodes’ neighborhoods to update their feature representations. This process results in the generation of enriched graph embeddings that capture both local and global semantic relations within the corpus.
A single convolutional layer of GCN can be calculated as:
$H=\widetilde{A}XW$ (3)
In this equation, $X\in\mathbb{R}^{n\times m}$ is the input matrix having ’n’ nodes with ’m’ dimensional features. $W\in\mathbb{R}^{m\times h}$ represents the weight matrix, and $\widetilde{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ defines the normalized symmetric adjacency matrix. Here $D$ is defined as $D_{ii}=\sum_{j}A_{ij}$. The $\widetilde{A}X$ term captures the part of the graph relevant to the input.
For a two-layer Vocabulary GCN with a ReLU activation function, the representation is as follows:
$VGCN=\text{ReLU}(X_{mv}\widetilde{A}_{vv}W_{vh})W_{hc}$ (4)
where ’m’ represents the mini-batch size, ’c’ is the size of the sentence embedding. ’v’ is the vocabulary size and ’h’ is the size of the hidden layer.
4.2.3 Combined GCN-BERT
A sentence containing entities is extracted from the abstract and fed to GCN-BERT. This sentence containing these target entities is processed as follows:
  • 1.
    Tokenization: We first tokenize each input sentence using the BioBERT tokenizer.
  • 1.
    Embeddings: The resulting tokenized sentence, containing a specific entity pair, is passed through the BioBERT embedding layer to generate embeddings for each token in the sentence.
  • 2.
    Vocabulary Graph Embeddings: Simultaneously, the tokenized input sentence is processed through the Graph Convolutional Network (GCN) to extract graph embeddings (as indicated by the blue ovals in Figure 2) for the relevant part of the input, which represent relationships between entities.
  • 3.
    Concatenation: The graph embeddings are concatenated with the BioBERT embeddings, resulting in an ensembled embedding representation.
  • 4.
    GCN-BERT: This combined embedding is then passed through the BERT model, which consists of a stack of 12 encoders in a feedforward neural network. This configuration is a standard choice in BERT models. The model is used for bidirectional training to produce the final vector representation.
  • 5.
    Classification: Finally, this vector is passed through a fully connected layer for the final classification task.
5 Experimental Setup
We explored a range of tuning techniques and combinations to optimize both the BERT and GCN-BERT model.
5.1 Entity Masking
Entity masking is a crucial technique for enhancing the model’s ability to generalize relationships from text. We employed two entity masking strategies:
  • 1.
    Masking Target Entities: Replacing the target entities with their entity type while leaving non-target entities unchanged, also known as reverse entity masking. This strategy resulted in the highest accuracy in our experiments.
  • 2.
    Masking Non-Target Entities: Masking the non-target entities from the sentence while keeping target entities intact.
5.2 Downsampling Negative Samples
The negative class, augmented with generated negative class instances as described in Section 3, ended up being much larger than the other classes. To address this class imbalance in our dataset, we implemented a 0.6 downsampling ratio for negative samples. This technique significantly improved the model’s ability to handle class imbalance.
5.3 Weighted Loss
We incorporated weighted loss during training to tackle class imbalance. Weighted loss assigns higher importance to minority classes, effectively addressing the imbalanced sample distribution. This technique was particularly effective given the prevalence of negative class samples in our dataset. The weighted loss can be calculated as:
$\text{Weighted Loss}=\sum_{i=1}^{N}w_{i}\cdot L(y_{i},f(x_{i}))$
where $N$ is the number of samples, $w_{i}$ is the weight assigned to the i-th sample and $L(y_{i},f(x_{i}))$ is the loss for the i-th sample, typically the cross-entropy loss.
6 Results
@@ 6.1 Evaluation Metrics
We assessed the model’s performance using several evaluation metrics, including precision, recall, Micro F1, Macro F1, and Weighted F1 scores.
The Macro F1 score represents the average of F1 scores across all classes and can be calculated as follows:
[…]
where $N$ is the number of classes, and $F1_{i}$ is the F1 score for class $i$.
The Weighted F1 score is computed by taking the mean of per-class F1 scores while considering the support of each class. In this context, ”support” refers to the proportion of each class’s representation relative to the total number of instances. It can be calculated as:
$\text{Weighted F1}=\frac{1}{N}\sum_{i=1}^{N}\frac{N_{i}}{N}\times F1_{i}$
where $N$ is the number of classes, $N_{i}$ is the number of true samples in class $i$, and $F1_{i}$ is the F1 score for class $i$.
The Micro F1 score, often referred to as accuracy, measures the proportion of correctly classified observations out of the total number of observations and can be calculated as:
$\text{Micro F1}=\frac{2\times\text{TP}}{2\times\text{TP}+\text{FP}+\text{FN}}$
6.2 Merged Dataset Analysis
We conducted a comprehensive comparison between the performance of our merged dataset and the ChemProt dataset processed by Sun et al. [35] For this evaluation, we specifically focused on CPR groups 3, 4, 9, and 10, as they are common to both datasets processed by Sun et al. The substantial increase in the size of our dataset led to a significant improvement in performance. The results for various CPR groups are presented in Table 3, with Precision, Recall and $F_{1}$ scores averaged over five execution runs. Notably, our model demonstrates a marked increase in performance across all positive groups (CPR 3, 4, and 9), attributed to the augmented sample counts within these groups. By concentrating on CPR groups 3, 4, and 9, shared between the ChemProt and DrugProt datasets, we aim to underscore the effectiveness of our approach. These groups serve as a crucial benchmark for evaluating the impact of our merged dataset, representing interactions consistently captured across both datasets. Analyzing the performance enhancements in these shared groups highlights how increased sample counts resulting from the merging of ChemProt and DrugProt contribute to the overall effectiveness of our BioBERT model. This focused comparison demonstrates the tangible benefits of our methodology in accurately predicting chemical-gene interactions.
ChemProt ChemProt-DrugProt
P R F P R F
CPR-3 0.7372 0.7407 0.7384 0.8690 0.8785 0.8735
CPR-4 0.7951 0.8236 0.8083 0.8894 0.9263 0.9073
CPR-9 0.6026 0.6976 0.6410 0.7620 0.8520 0.8040
CPR-10 0.9544 0.9415 0.9479 0.9725 0.9575 0.9649
Micro F1 0.7839 0.8875
Std dev 0.0025 0.0021
The BioBERT [16] model offers two configurations: ’Base’ and ’Large.’ BioBERT-Large’s pre-training was performed on a more extensive vocabulary that encompasses a broader spectrum of biomedical terms compared to BioBERT Base. BioBERT-Base was used for the experiments in Table 3. To comprehensively assess performance on our merged dataset, we also conducted experiments with a BioBERT-Large model. Specifically, we developed three variants of the BioBERT approach: BioBERT-Base with the top model, BioBERT-Large without the top model, and BioBERT-Large with the top model.
The $F_{1}$ scores obtained from the test set across five individual runs are presented in Table 4. These results consistently demonstrate a pattern: BioBERT-Large consistently outperforms BioBERT-Base in terms of $F_{1}$ scores. Additionally, the inclusion of the top model leads to an enhancement in the $F_{1}$ score. However, note that this enhancement is not statistically significant, indicating that achieving state-of-the-art results does not necessarily require the adoption of a more complex system.
@@ 6.2 Merged Dataset Analysis
6.3 BioBERT and GCN-BERT Comparison
In this section, we compare the performance of incorporating GCN with BioBERT. To ensure robustness of our results, we report them as an average of five runs. The parameters and evaluation metrics are presented in Table 5. Table 6 presents results per class for Experiment 5 from Table 5 and compares the results with those obtained using BioBERT-Base and BioBERT-Large as described in Section 4.1. As noted in Section 3, CPR 7 and CPR 8 are not included in the results due to the low number of training data instances for those relations. The results show that incorporating the GCN does increase the overall F-1 score for some instances.
Table 5: GCN-BERT Experimental Results
Ex. # Parameters Micro F1 (std dev) Weighted avg F1 (std dev) Macro F1 (std dev)
1 Downsampling (ratio = 0.6) 83.253 (0.007) 83.313 (0.021) 71.2 (0.005)
2 No downsampling 87.348 (0.008) 87.184 (0.003) 69.2 (0.003)
3 Reverse entity masking without downsampling 89.146 (0.023) 89.057 (0.090) 74.69 (0.004)
4 BioBERT-Large without downsampling 88.618 (0.010) 88.453 (0.003) 71.78 (0.006)
5 Reverse entity masking with Downsampling (ratio 0.6) 89.326 (0.012) 89.438 (0.017) 75.68 (0.002)
Table 6: Comparing BioBERT-Base, BioBERT-Large and GCN-BERT
BioBERT-Base BioBERT-Large GCN-BERT
P R F P R F P R F
CPR-1 0.7899 0.8607 0.8233 0.8623 0.8871 0.8733 0.6929 0.6730 0.6874
CPR-2 0.5574 0.4903 0.5202 0.5380 0.5066 0.5213 0.7514 0.6298 0.6766
CPR-3 0.8333 0.8677 0.8494 0.8671 0.8809 0.8735 0.7419 0.7657 0.7535
CPR-4 0.8697 0.9253 0.8965 0.8929 0.9311 0.9116 0.8043 0.8322 0.8141
CPR-5 0.8898 0.8956 0.8920 0.8861 0.9319 0.9080 0.7238 0.7103 0.7128
CPR-6 0.8834 0.9402 0.9105 0.9276 0.9440 0.9357 0.7378 0.8087 0.7661
CPR-9 0.7325 0.8279 0.7772 0.8203 0.8157 0.8174 0.7372 0.6761 0.7053
CPR-10 0.9292 0.9140 0.9215 0.9319 0.9270 0.9294 0.9326 0.9395 0.9355
Micro F1 0.8836 0.8952 0.8933
Following the analysis of GCN-BERT’s results, a plausible pattern can be detected in its successes and shortcomings, particularly concerning the dominant relations that entities hold within a paragraph. Consider the following text:
Some clinically used compounds, such as acetazolamide, methazolamide, ethoxzolamide, dichlorophenamide, dorzolamide, brinzolamide, topiramate, sulpiride, and indisulam, or the orphan drug benzolamide, showed effective hCA VI inhibitory activity, with inhibition constants of 0.8-79 nM. The best inhibitors were brinzolamide and sulpiride (KI values of 0.8-0.9 nM), the latter compound being also a CA VI-selective inhibitor.
Here, the entity ”hCA VI” is related to most of the other entities in the paragraph, namely, dichlorophenamide, dorzolamide, brinzolamide, and topiramate, under the ”CPR-4” relation. The model is able to correctly classify these relations since ”CPR-4” is well-represented in the training data. However, the model encounters difficulty in cases where ”hCA VI” has a relation other than ”CPR-4”. Additionally, since the dataset is imbalanced, with many samples having a ”NOT” relation, there are many false negatives in the model’s predictions.
In summary, the performance of GCN-BERT appears to hinge on the frequency and dominance of specific relations associated with entities in a given paragraph. While the model excels in cases where an entity has the same type of relation with other entities, it faces challenges when confronted with less frequent or alternative relations, leading to misclassifications in certain instances.
7 Discussion
In this section, we discuss the our analysis of the results reported in Section 6.
7.1 Effect of Class Sample Size
In our analysis of the experiment results, we observed an expected relationship between the number of training samples in a class and the model’s performance. Specifically, the model consistently performed better on classes that had a higher number of samples in the training data.
For instance, the ’NO RELATIONS’ class (CPR-10) consistently had the highest number of samples, and as a result, the model consistently delivered the best performance on this class. This trend was also observed in the case of CPR-4, which had the second-highest number of samples and consistently yielded the second-best performance across different tuning techniques.
In Experiments 1 and 2 in Table 5, we compared the results of our model with and without downsampling of the negative samples. We found that, without downsampling in Experiment 2, the model achieved a higher micro F1 score compared to Experiment 1, which involved downsampling. The absence of downsampling in Experiment 2 led to more samples in the CPR-10 class, resulting in better performance in this specific class. However, due to the imbalance in the dataset, Experiment 2 performed worse than Experiment 1 in other classes. This discrepancy is evident when looking at the macro F1 score, which reflects the overall performance across all classes.
7.2 Handling class imbalance
Given the presence of an imbalanced dataset with numerous negative samples, we employed specific techniques to address this challenge, including downsampling and weighted loss. Notably, our findings indicate that a downsampling ratio of 0.6 offers an optimal balance for handling the data imbalance. This strategy not only aids in mitigating class imbalance issues but also enables the model to allocate sufficient attention to various relation classes beyond the negative samples.
7.3 Entity Masking
Our experiments confirm the effectiveness of masking target entities with their corresponding semantic types, leading to a significant 6% increase in accuracy compared to the alternative approach. This result implies that the presence of other entities within the sentence plays a substantial role in enhancing the model’s contextual understanding, thereby improving its ability to predict relations among entities.
7.4 Graph Embeddings
While the vocabulary graph embeddings contributes to the GCN-BERT’s ability to capture the global context, it is worth noting that the vocabulary graph is constructed using all words within the corpus. This approach, though informative, may introduce a potential challenge for the model. The graph’s inclusion of nodes for all words in the corpus provides an abundance of information, which might inadvertently introduce noise, diverting the model’s focus from its primary task of capturing relationships between chemical entities. One possible improvement could involve constructing the vocabulary graph using only entity words.
7.5 Difficulty handling unseen words during inference
One potential drawback of the GCN-BERT approach is its reliance on a vocabulary graph constructed using words from both the training and test corpora. This design choice might hinder the model’s performance on unseen data containing words that were not part of the original corpus. Specifically, generating graph embeddings for such words during inference can be challenging because nodes corresponding to those words are absent from the vocabulary graph.
8 Future Work
First, given the performance gains observed from merging the ChemProt and DrugProt datasets, we plan to explore additional data sources and methods for combining them to create even larger, more extensive datasets across a more diverse set of relationship types.
Second, based on our observations, we propose that the below avenues will help further enhance the performance of both the BioBERT and GCN-BERT models. Firstly, better handling of OOV words. The vocabulary graph contains nodes, which are word tokens tokenized by the BioBERT tokenizer. It splits OOV words into subwords and uses ’##’ to indicate that they are part of a larger word. For example, the word ’xyz’ may be tokenized as ’x’ and ’##yz.’ This issue is extremely apparent with chemical tokens. We believe that incorporating even more domain-specific knowledge into the text corpus to generate these embeddings and tokens could enhance the representation of chemical and gene entities within the vocabulary graph. An example of such an approach is [44], which uses a subword embedding model and incorporates domain knowledge from MeSH. Specifically for GCN-BERT, Since the vocabulary graph is constructed using word tokens as nodes, the adoption of a domain-specific tokenization scheme or the inclusion of vocabulary words in the text corpus could lead to more coherent graph embeddings.
Third, for GCN-BERT utilizing a filtered vocabulary graph. The vocabulary graph is typically constructed using all words in the vocabulary, which can introduce noise into the model. To address this, we plan to implement a ”Filtered Vocabulary Graph” that contains only entities as nodes. This approach is inspired by the work of [42], who constructed a mention graph that included both entity mention and document nodes and applied Graph Convolutional Networks to transform it into an entity graph. Focusing more on entities in the vocabulary graph could yield more accurate and informative results for our model.
Lastly, separate the training of the GCN model and BERT. As suggested by previous authors [24], we plan to explore the possibility of training the Graph Convolutional Network and BERT separately. Subsequently, we will combine the embeddings before passing them through the final classification layer. Training them separately might enhance the learning capabilities of both models, leading to improved representations and, consequently, enhanced model performance.
9 Conclusion
This paper presents a new, larger, more comprehensive data set created by merging the ChemProt and DrugProt datasets to augment sample counts and improve model accuracy. We evaluated the merged data set using two relationship extraction models: BioBERT and GCN-BERT. The BioBERT architecture aims to capture contextualized information contained within the sentence, while the GCN-BERT architecture aims to classify relations by leveraging both global information through graph embeddings and local contextual information through BERT. We demonstrated significant improvements in model performance, particularly in CPR groups shared between the datasets.
10 Data and Software Availability
The data and software are publicly available at: https://github.com/bmcinnes/UCSD-VCU-IE-Collaboration
Acknowledgement
This work was funded in part by NSF award numbers 1939885 and 1939951, and utilized computer resources provided in part by NSF award numbers 1730158, 2100237, 2120019 and 2316003.